OpenStudy (anonymous):

find k give that |1 k 3 | |k 1 -1| = 7 |3 4 2 | btw.. this is suposed to be a matix

OpenStudy (anonymous):

woops matrix

OpenStudy (anonymous):

it is clear what you have to do right? take the determinant and set it equal to 7, solve for k

OpenStudy (anonymous):

that is, solve \[-2 k^2+9 k-3=7\] a quadratic equation

OpenStudy (anonymous):

yahh but like my result doesnt make any sence

OpenStudy (anonymous):

yas i got that equation.. but like my result does make sence

OpenStudy (jamesj):

Why not? What values of k did you find?

OpenStudy (anonymous):

like 1.113406013 and then -5.6134

OpenStudy (anonymous):

that makes no sence of me ha

OpenStudy (jamesj):

no, that's right. \[ 2k^2 - 9k + 10 = 0 \] Now use the quadratic formula.

OpenStudy (anonymous):

its right? yah i used that one

OpenStudy (anonymous):

i checked in the answer booklet it says like 2 or 5/2

OpenStudy (anonymous):

Yes, after applying Laplace expansion to that matrix you will get a quadratic equation equation which will be \( 2k^2 - 9k + 10 = 0 \) and now if you solve for k then you will get k=2 or \(K = \frac 5 2 \).

OpenStudy (anonymous):

now how do you get that though?

OpenStudy (anonymous):

Factor \[ 2k^2 - 9k + 10 = 0 \]

OpenStudy (jamesj):

c'mon. In standard notation a = 2, b = -9, c = 10. Hence \[ b^2 - 4ac = 81 - 80 = 1 \] and therefore the roots \[ k = \frac{ -b \pm \sqrt{b^2-4ac}}{2a} = \frac{9 \pm 1}{4} = 2 \ or \ 5/2 \]

OpenStudy (anonymous):

`thank you!!!